Suppose you have a population of flour beetles with 1,000 individuals. Normally the beetles are red; however, this population is polymorphic for a mutant autosomal body color, black, designated by bb. Red is dominant to black, so BB and Bb genotypes are red. Assume the population is at Hardy–Weinberg equilibrium, with equal frequencies of the two alleles. What would be the expected frequencies of the red and black phenotypes?

Select one:

a. All red

b. 0.5 red; 0.5 black

c. All black

d. 0.75 red; 0.25 black

e. 0.25 red; 0.75 black
Answer #1

Hello

According to hardy-weinberg law relationship between gene frequency and genotype frequency can be expreessed as

p2+2pq+q2 =1

Or (p+q)2 = 1

Or p+q= 1

According to question population is in hardy weinberg equilibrium, with equal frequencies of two alleles .Therefore genotypic frequency of alleles is

p = 0.5

q= 0.5

number of beetles have Red colour is ( BB+Bb)

BB = 0.5×0.5×1000 = 250

and   Bb= 2×0.5×0.5×1000=500  

Therefore    BB+ Bb = 750

number of beetles have  black colour is bb

   = 0.5×0.5×1000=250

Therefore phenotypic frequency is red :black

   =750: 250 or 0.75 red;0.25 black

   Therefore answer is d.

note ; if you satisfied with answer. Please give postive feedback